Four Times the sum of three consecutive integers is 48.
a. write an equation that could be used to find the integers
b. what are the integers
I believe the answer is Integers: 3,4, and 5 because 3+4+5 = 12 * 4 = 48. but I do not know how to write the equation.

Answer by ankor@dixie-net.com(22740)

Four Times the sum of three consecutive integers is 48.
a. write an equation that could be used to find the integers
b. what are the integers
I believe the answer is Integers: 3,4, and 5 because 3+4+5 = 12 * 4 = 48. but I do not know how to write the equation.
:
4(x + (x+1) + (x+2)) = 48
4(3x + 3) = 48
12x + 12 = 48
12x = 48 - 12
12x = 36
x = 3